# Spineless Traversal for Layout Invalidation

### v4

@@ 7.4. Case Study: Twitter
We now focus specifically on our trace of Twitter (now X), a social media platform. This trace of 125 frames captures the user opening the Twitter news feed, loading the default number of tweets, and scroll down repeatedly to load more tweets. Twitter is a large web page, and the tree grows to 3 700 DOM nodes with a depth of 53 and a fanout of 128. Considering all the frames in aggregate, Twitter sees a mean speedup of 1.99×1.$9$ over the Double Dirty Bit algorithm.
Most of the 125 incremental layouts are small, dirtying no more than 20 nodes (Figure 16). For these frames, incremental layout spends most of its time accessing auxiliary nodes. However, the largest incremental layout dirties several hundred nodes. We now discuss common kinds of frames.
@@
Twitter uses a lazy-loading technique which first loads a “shell” page and then gradually adds more and more elements to the shell as more content is loaded over the network. For example, the header bar, side bar, ads, and tweets all load separately and require separate incremental layouts. Scrolling causes yet more content (tweets and ads) to load. Each of these frames typically involve inserting a single large subtree. Allocating the new nodes’ OM objects (and possibly rebalancing the OM data structure) makes these frames difficult for Spineless Traversal despite its bulk insertion optimizations, resulting in a slowdown over Double Dirty Bit, typically by about 2×2$\times$2 ×. That said, the latency is partially hidden by the network latency of loading the content in the first place, so the slowdown here may be less critical than for other frames.
@@
Moreover, some frames mix file loading, lazy loading, and removals, probably at the whim of the task scheduler. Often many images load in at once. In this case the time taken for Spineless Traversal basically sums over the time for each individual change, while Double Dirty Bit can amortize the cost of traversing auxiliary nodes. These frames are often small, so Spineless Traversal’s speed-ups are substantial, but probably smaller than if each modification was laid out in its own frame.
@@ 8. Related Work
Speeding up computations by reusing previously-computed results is a long-studied topic in computer science broadly (Michie, 1968) and programming language theory in particular; Liu (2024) and Ramalingam and Reps (1993) give thorough surveys of the field. The recent Self-Adjusting Computation (SAC) (Acar, 2005) framework proposes incrementalizing arbitrary computations, including a cost semantics (Acar, 2009), optimizations for data structure operations (Acar et al., 2010), and opportunities for parallelization (Anderson et al., 2021). The Adapton framework (Hammer et al., 2014) aims at demand-driven incremental computation, and allows manually-specified annotations Hammer et al. (2015) for greater reuse. While this prior work focuses on general-purpose computations, Spineless Traversal is focused on a particularly critical application: web browser layout. This application-specific focus has precedent: Pugh and Teitelbaum (1989) speed up memoization for functional programs over lists using “chunky decomposition”, while differential dataflow databases (McSherry et al., 2013) that incrementalize relational algebra are prominent in industry.
@@
Reps (1982) wrote the earliest work on incremental evaluation of attribute grammars, motivated by syntax-directed editors; later work (Reps et al., 1986) allows references to non-neighbor attributes. However, these early papers require recomputation immediately after every tree change, whereas web browsers, our target application, batch multiple updates and perform layout only once per frame. The standard in web browsers is instead the Double Dirty Bit algorithm, implemented in all existing web rendering engines. Industry publication web.dev ran an article on Double Dirty Bit (Garseil, [n. d.]), and the algorithm also features in textbooks (Panchekha and Harrelson, 2024).
@@
The formal methods community has put significant effort into formalizing web page layout. Jones et al. ([n. d.]) proposed using attribute grammars, similar to the DSL in Figure 3, for formalizing web-like layout rules. Later work (Chen et al., 2022) proposes synthesizing schedules from the attribute grammar rules, including proposals (Meyerovich and Bodik, 2010; Meyerovich, 2013) to use parallel schedules to further improve layout performance. The Cassius project (Panchekha and Torlak, 2016) formalized a significant fragment of CSS 2.1 using an attribute-grammar-like formalism. Our layout implementation is based on Cassius. Later work (Panchekha et al., 2018) also proposed using the Cassius formalism to verify web page layouts, including in a custom proof assistant (Prazina et al., 2023). However, none of these works investigate incremental layout. By contrast, the Medea project (Liu et al., 2023) proposed synthesizing incremental layout algorithms by automatically synthesizing dirty bit propagation code. Our work extends Medea by exploring optimized incremental traversal algorithms.

### v5

@@ 7.4. Case Study: Twitter
We now focus specifically on our trace of Twitter (now X), a social media platform. This trace of 125 frames captures the user opening the Twitter news feed, loading the default number of tweets, and scrolling down repeatedly to load more tweets. Twitter is a large web page, and the tree grows to 3 700 DOM nodes with a depth of 53 and a fanout of 128. Considering all the frames in aggregate, Twitter sees a geometric mean speedup of 1.99×1.$9$ over the Double Dirty Bit algorithm.
Most of the 125 incremental layouts are small, dirtying no more than 20 nodes (Figure 16). For these frames, Double Dirty Bit spends most of its time accessing auxiliary nodes. However, the largest incremental layout dirties several hundred nodes. We now discuss several common kinds of frames in the Twitter trace.
@@
Twitter uses a lazy-loading technique which first loads a “shell” page and then gradually adds more and more elements to the shell as more content is loaded over the network. For example, the header bar, side bar, ads, and tweets all load separately and require separate incremental layouts. Scrolling causes yet more content (tweets and ads) to load. Each of these frames typically involve inserting a single large subtree. Allocating the new nodes’ OM objects (and possibly rebalancing the OM data structure) makes these frames difficult for Spineless Traversal despite its bulk insertion optimizations, Spineless Traversal is slower than Double Dirty Bit, typically by about 2×2$\times$2 ×. That said, the latency is partially hidden by the network latency of loading the content in the first place, so the slowdown here may be less critical than for other frames.
@@
Moreover, some frames mix file loading, lazy loading, and removals, probably at the whim of the task scheduler. Often many images load in at once. In this case the time taken for Spineless Traversal basically sums over the time for each individual change, while Double Dirty Bit can amortize the cost of traversing auxiliary nodes. These frames are often small, so Spineless Traversal’s speed-ups are still substantial, but probably smaller than if each modification was laid out in its own frame.
@@ 8. Related Work
Speeding up computations by reusing previously-computed results is a long-studied topic in computer science broadly (Michie, 1968) and programming language theory in particular; Liu (2024) and Ramalingam and Reps (1993) give thorough surveys of the field. The recent Self-Adjusting Computation (SAC) (Acar, 2005) framework proposes incrementalizing arbitrary computations, including a cost semantics (Acar, 2009), optimizations for data structure operations (Acar et al., 2010), and opportunities for parallelization (Anderson et al., 2021). The Adapton framework (Hammer et al., 2014) aims at demand-driven incremental computation, and allows manually-specified annotations (Hammer et al., 2015) for greater reuse. While this prior work focuses on general-purpose computations, Spineless Traversal is focused on a particularly critical application: web browser layout. This application-specific focus has precedent: Pugh and Teitelbaum (1989) speed up memoization for functional programs over lists using “chunky decomposition”, while differential dataflow (McSherry et al., 2013), which incrementalizes relational algebra in databases, is prominent in industry.
@@
Reps (1982) wrote the earliest work on incremental evaluation of attribute grammars, motivated by syntax-directed editors; later work (Reps et al., 1986) allows references to non-neighbor attributes. However, these early papers require recomputation immediately after every tree change, whereas web browsers, our target application, batch multiple updates and perform layout only once per frame. The standard in web browsers is instead the Double Dirty Bit algorithm, described in industry publications (Garseil, [n. d.]) and textbooks (Panchekha and Harrelson, 2024).
@@
The formal methods community has put significant effort into formalizing web page layout. Jones et al. ([n. d.]) proposed using attribute grammars, similar to the DSL in Figure 3, for formalizing web-like layout rules. Later work (Chen et al., 2022) proposes synthesizing schedules from the attribute grammar rules, including proposals (Meyerovich and Bodik, 2010; Meyerovich, 2013) to use parallel schedules to further improve layout performance, though these proposals have not proven practical (servo-no-parallel). The Cassius project (Panchekha and Torlak, 2016) formalizes a significant fragment of CSS 2.1 using an attribute-grammar-like formalism. Our layout implementation is based on Cassius. Later work also proposed using the Cassius formalism to verify web page layouts (Panchekha et al., 2018), including in a custom proof assistant (Prazina et al., 2023). However, none of these works investigate incremental layout. By contrast, the Medea project (Liu et al., 2023) proposed synthesizing incremental layout algorithms by automatically synthesizing dirty bit propagation code. Our work extends Medea by exploring optimized incremental traversal algorithms.
Acknowledgements.
We thank Andreas Kling and Chris Harrelson for many conversations discussing implementation challenges in layout invalidation and Ian Kilpatrick for his discussion of reflow roots. Ryan Stutsman and Anton Burtsev for helped with measuring low-level performance. We also thank the anonymous reviewers for their many comments and suggestions, which greatly improved the paper. This work was funded by the National Science Foundation under Grant No. CCF-2340192 and by a Google Faculty Research Award.